Superior Markets, Inc., operates three stores in a large metropolitan area. A segmented absorption costing income statement for the company for the last quarter is given below:

Superior Markets, Inc.
Income Statement
For the Quarter Ended September 30
Total North
Store
South
Store
East
Store
Sales $ 4,200,000 $ 840,000 $ 1,680,000 $ 1,680,000
Cost of goods sold 2,310,000 500,000 886,000 924,000
Gross margin 1,890,000 340,000 794,000 756,000
Selling and administrative expenses:
Selling expenses 841,000 243,400 321,000 276,600
Administrative expenses 443,000 118,000 168,900 156,100
Total expenses 1,284,000 361,400 489,900 432,700
Net operating income (loss) $ 606,000 $ (21,400 ) $ 304,100 $ 323,300

The North Store has consistently shown losses over the past two years. For this reason, management is giving consideration to closing the store. The company has asked you to make a recommendation as to whether the store should be closed or kept open. The following additional information is available for your use:

The breakdown of the selling and administrative expenses that are shown above is as follows:

Total North
Store
South
Store
East
Store
Selling expenses:
Sales salaries $ 256,800 $ 68,600 $ 78,200 $ 110,000
Direct advertising 177,000 63,000 84,000 30,000
General advertising* 63,000 12,600 25,200 25,200
Store rent 285,000 81,000 114,000 90,000
Depreciation of store fixtures 22,000 5,800 7,200 9,000
Delivery salaries 24,600 8,200 8,200 8,200
Depreciation of delivery
equipment
12,600 4,200 4,200 4,200
Total selling expenses $ 841,000 $ 243,400 $ 321,000 $ 276,600

*Allocated on the basis of sales dollars.

Total North
Store
South
Store
East
Store
Administrative expenses:
Store managers' salaries $ 88,000 $ 27,000 $ 36,000 $ 25,000
General office salaries* 63,000 12,600 25,200 25,200
Insurance on fixtures and inventory 37,000 11,100 15,000 10,900
Utilities 85,140 28,840 28,560 27,740
Employment taxes 64,860 17,460 22,140 25,260
General office—other* 105,000 21,000 42,000 42,000
Total administrative expenses $ 443,000 $ 118,000 $ 168,900 $ 156,100

*Allocated on the basis of sales dollars.

The lease on the building housing the North Store can be broken with no penalty.

The fixtures being used in the North Store would be transferred to the other two stores if the North Store were closed.

The general manager of the North Store would be retained and transferred to another position in the company if the North Store were closed. She would be filling a position that would otherwise be filled by hiring a new employee at a salary of $11,600 per quarter. The general manager of the North Store would continue to earn her normal salary of $12,600 per quarter. All other managers and employees in the North store would be discharged.

The company has one delivery crew that serves all three stores. One delivery person could be discharged if the North Store were closed. This person’s salary is $5,200 per quarter. The delivery equipment would be distributed to the other stores. The equipment does not wear out through use, but does eventually become obsolete.

The company pays employment taxes equal to 15% of their employees' salaries.

One-third of the insurance in the North Store is on the store’s fixtures.

The “General office salaries” and “General office—other” relate to the overall management of Superior Markets, Inc. If the North Store were closed, one person in the general office could be discharged because of the decrease in overall workload. This person’s compensation is $6,300 per quarter.

Required:

1. How much employee salaries will the company avoid if it closes the North Store?

2. How much employment taxes will the company avoid if it closes the North Store?

3. What is the financial advantage (disadvantage) of closing the North Store?

4. Assume that the North Store's floor space can’t be subleased. However, let's introduce three more assumptions. First, assume that if the North Store were closed, one-fourth of its sales would transfer to the East Store, due to strong customer loyalty to Superior Markets. Second, assume that the East Store has enough capacity to handle the increased sales that would arise from closing the North Store. Third, assume that the increased sales in the East Store would yield the same gross margin as a percentage of sales as present sales in the East store. Given these new assumptions, what is the financial advantage (disadvantage) of closing the North Store?
Answer #1
1 Salaries which can be avoided,
Sales salaries 68600
Delivery salaries 5200
Store management salaries (27000-12600) 14400
Salary of new manager 11600
General office compensation 6300
Total 106100
2 Employement taxes which can be avoided=Salaries which can be avoided*15%=106100*15%=$15915
3 Gross margin lost if north store is closed -340000
Savings in cost:
Store rent 81000
Sales salaries 68600
Direct adveritising 63000
Delivery salaries 5200
Store management salaries 14400
Salary of new manager 11600
General office compensation 6300
Insurance (11100*2/3) 7400
Utilities 28840
Employment taxes 15915 302255
Decrease in net profit if North store is closed -37745
Financial disadvantage of closing North store=$37745
4 Gross margin earned from East store
Gross margin % of east store=Gross margin/Sales=756000/1680000=0.45=45%
Sales transferred from North to East store=1/4 th of sales=840000*1/4=$ 210000
Gross margin earned from sales transferred=210000*45%=$ 94500
If floor space can't be subleased,Then threre is no savings in store rent.
Savings in cost=302255-81000=$221255
Gross margin lost if north store is closed -340000
Gross margin earned from East store (Note:2) 94500
Net gross margin lost -245500
Savings in cost 221255
Net disadvantage of closing north store -24245